A company will sell Thingamabobs to consumers at a price of $106 per unit. The variable cost to produce Thingamabobs is $42 per unit. The company expects to sell 13,000 Thingamabobs to consumers each year. The fixed costs incurred each year will be $180,000. There is an initial investment to produce the goods of $3,200,000 which will be depreciated straight line over the 15 year life of the investment to a salvage value of $0. The opportunity cost of capital is 8% and the tax rate is 37%.

What is operating cash flow each year?  
Answer #1

Sales = 106 * 13,000 = $1,378,000

Variable costs = 13,000 * 42 = $546,000

Fixed costs = $180,000

Depreciation = 3,200,000/15 = $213,333.333333333

EBIT = Sales - Variable costs - Fixed costs - Depreciation

EBIT = 1,378,000 - 546,000 - 180,000 - 213,333.333333333

EBIT = $438,666.666667

Net Income = EBIT * (1 - tax rate)

Net Income = 438,666.666667 * (1 - 0.37)

Net Income = $276,360.00000021

Operating cash flow = Net Income + Depreciation

Operating cash flow = 276,360.00000021 + 213,333.333333333

Operating cash flow = $489,693.333333543

Can you please upvote? Thank You :-)